Potential energy of a satellite having mass m and rotating at a height of 6.4×10`6 m from the earth surface is ?

Grade:12

1 Answers

Khimraj
5 years ago
Potential energy 
PE= -GMm/(R+h)
G = 6.674 08 x 10-11 m3 kg-1 s-2
M = 5.9722×1024 kg
m = mass of satellite
R = 6.371*106 m
h = 6.4*106 m
Hope it clears.